# Looking for the $\gamma$-Ray Cascades of the KM3-230213A Neutrino Source

### v1

@@ 1 Introduction
The unexpected detection of a signal by the KM3NeT/ARCA Collaboration that is consistent with a $\gtrsim$100 PeV neutrino has potentially opened a new window into the most extreme accelerators in the Universe [1]. If confirmed to be of astrophysical origin, KM3-230213A stands out as the highest energy neutrino observed to date, raising questions regarding its astrophysical [2, 3, 4, 5] or cosmological [6] origins, as well as potential multiwavelength counterparts. This event prompted a variety of theoretical interpretations involving beyond Standard Model physics [7], including Lorentz invariance violation [8, 9, 10] and dark matter decay or interactions [11, 12]. Scenarios involving evaporating primordial black holes have also been explored [13]. The tension between the KM3NeT signal and the lack of similar detections by the larger IceCube detector motivates scenarios where the KM3NeT signal is produced by a transient source fortuitously located in the optimal region for KM3NeT detection [14, 15].
If the neutrino flux is generated via the hadronic interactions of ultra-high energy protons or nuclei, there should also be an accompanying flux of high-energy $\gamma$-rays [16, 17, 18, 19]. These $\gamma$-rays can initiate electromagnetic cascades on the extragalactic background light (EBL), shifting the bulk of their energy into the GeV–TeV band, where they could be observable by the Fermi Large Area Telescope (LAT). In the case of steady-state sources, these cascades could produce a diffuse $\gamma$-ray halo surrounding the source, the size and spectrum of which would depend sensitively on the intergalactic magnetic field (IGMF) strength. In the case of transient sources, the spectral evolution of the cascade is more difficult to model, however, the spatial evolution is simpler, as the $\gamma$-ray cascade produced over short time delays has a spatial extension much smaller than the resolution of current $\gamma$-ray instruments.
@@ 2.1 Electromagnetic Cascade Modeling
[…]
@@
where $\eta\sim\mathcal{O}(1)$ accounts for the uncertainty in the production mechanism (e.g. $\eta=1$ for photo-pion production or $\eta=1/2$ for proton-proton interactions), and $dN(z)/dE_{\nu}d_{A}=4\pi\Phi_{\nu}\tau_{\text{ARCA}}$ is the integrated differential number flux of neutrinos detected over an active observation time of KM3NeT/ARCA high-energy neutrino detector, $\tau_{\text{ARCA}}\approx 1\;\text{year}$. Furthermore, we assume that the $\gamma$-rays are produced with a spectral shape that is centered about 440 PeV and whose width is narrow enough such that essentially all the injected energy is reprocessed into cascade photons. Practically, we model this as a log-normal distribution with a variance (width) of up to a decade. We note that our results are not sensitive to the spectral width of the injected $\gamma$-rays as long as this width is less than $\sim$3 decades. At larger widths, there would be a significant number of unattenuated TeV $\gamma$-rays for distances below $\sim$10 Mpc. If the emission extends below 100 GeV, there would be an additional unattenuated $\gamma$-ray component which may be bright compared to the KM3NeT flux, and would typically strengthen the resulting constraints in this paper.
@@
All the primary $\gamma$-rays in our model will be attenuated due to pair production with the background radiation field stemming from a combination of the cosmic microwave background (CMB) and the EBL. We adopt the EBL model from Saldana-Lopez et al. (2021) [22]. After $e^{+/-}$ (hereafter, collectively called electrons) pair produce, these electrons are able to inverse-Compton scatter with the same background radiation, producing $\gamma$-rays. This cascade cycle reprocesses the energy above $\sim\mathcal{O}(\text{TeV})$ into GeV photons that are visible to the Fermi-LAT. We compute these cascades semi-analytically using $\gamma$-Cascade [20].
@@
As the primary $\gamma$-rays leave the source at comoving distance $d_{s}$, they will cascade and change their original direction of travel. This is caused by the motion of the electrons under the influence of the IGMF. Once electrons are generated via pair production, they experience an acceleration perpendicular to their velocity, which causes them to turn with a Larmor radius, $R_{L}=E_{e}/eB(z)$, where $B(z)=B_{0}(1+z)^{-\alpha}$ is the IGMF strength at redshift $z$. Here we take $\alpha=0$ for a non-evolving IGMF and $\alpha=4$ for IGMF evolution driven by a dynamo mechanism. The total angle by which the electrons’ trajectory becomes deflected is given by,
[…]
@@
and where $u_{\text{CMB}}$ is the energy density of the background radiation field, $\sigma_{T}$ is the Thompson cross section, and $\gamma_{e}=E_{e}/m_{e}c^{2}$ is the gamma factor of the electrons. We can evaluate the integral above in terms of the energy-loss distance of the electrons, $d_{\text{ICS}}=cE_{e}/(dE_{e}/dt)$,
[…]
$=\frac{3m_{e}^{2}c^{4}eB(z)}{4\sigma_{T}u_{\text{CMB}}(z)E_{e}^{2}}$
$=\frac{3m_{e}^{2}c^{4}eB_{0}}{4\sigma_{T}u_{\text{CMB},0}E_{e}^{2}(1+z)^{4+\alpha}}.$ (2.4)
@@
We further assume a small-angle approximation, where $\delta\ll\pi$, (i.e. $d_{\text{ICS}}\ll R_{L}$). As the cascade progresses, the mean free path of the photons grows rapidly below $E_{\gamma}\sim\mathcal{O}(\text{PeV})$. At energies above the point where the mean free path reaches its minimum, the mean free path scales as $D_{\gamma\gamma}=\langle\bar{\sigma}_{\gamma\gamma}(E_{\gamma})n_{\text{CMB}}\rangle^{-1}\propto E_{\gamma}$. Therefore, one can think of the cascade as being composed of several cycles that extend out from the source up to characteristic distances set by the mean free path of the second-to-last pair-production step, $d_{c}\approx D_{\gamma\gamma}(E_{\gamma}/2)$. After this point, the original direction of the cascading photons has been changed by $\delta$ from it’s original trajectory.
@@
The “last” photon is the one that is detected on Earth by the observer coming from a direction $\theta_{\text{obs}}$ with respect to the line of sight to the source. It’s been shown in several studies that this angle, which manifests in the characteristic apparent angular extension of the source, can be approximated by the following,
$\theta_{\text{obs}}(E_{\gamma},z)$ $\approx\frac{D_{\gamma\gamma}(E_{\gamma}^{0}/2,z)}{d_{A}}\delta(E_{e}=E_{\gamma}^{0},z)$
$\approx(1+z)^{-(4+\alpha)}\theta_{\text{obs}}(E_{\gamma},z=0),\text{with}$ (2.5)
@@
$\theta_{\text{obs}}(E_{\gamma},z=0)$ $\approx 1^{\prime\prime}\times\left(\frac{E_{\gamma}}{100\;\text{GeV}}\right)^{-1}\left(\frac{B_{0}}{10^{-14}\;\text{G}}\right),$ (2.7)
@@
where $d_{A}=d_{e}/(1+z)$ is the angular diameter distance. We approximate the profile of the photons arriving at earth by a Gaussian distribution $\mathcal{N}(\mu_{\theta}=0,\sigma_{\theta}=\theta_{\text{obs}})$ centered at $\mu_{\theta}=0$ with a characteristic width of $\sigma_{\theta}=\theta_{\text{obs}}(E_{\gamma},z)$. For a given observed angle, the observed time delay due to the additional travel distance of the cascade is given by,
$\tau_{\text{obs}}(E_{\gamma},z)$ $\approx(1+z)\frac{D_{\gamma\gamma}(E_{\gamma}^{0},z)}{2}\delta^{2}(E_{\gamma}^{0},z)$
@@
$\approx 5\,\text{yrs}\times(1+z)^{-(8+\alpha)}\left(\frac{D_{\gamma\gamma}}{10\;\text{Mpc}}\right)\left(\frac{B_{0}}{10^{-14}\;\text{G}}\right)^{2}\left(\frac{E_{\gamma}}{100\;\text{GeV}}\right)^{2}.$ (2.8)
@@
Notice that since $\tau_{\text{obs}}\propto\theta_{\text{obs}}^{2}$, the time delay of the $\gamma$-rays is distributed as a chi-squared distribution with one degree of freedom, $\chi^{2}_{0}((\tau-\tau_{s})/\tau_{\text{obs}})$, where $\tau_{s}$ is the light-travel time to the source and $\tau$ is the time at which an observation is made . This is a consequence of our simplified model where the photons’ arrival is distributed as an angular Gaussian. A key result is that the total integrated flux of $\gamma$-rays collected after a total observation time $T_{\text{obs}}$ is given by the CDF of this chi-squared distribution,
$E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},T_{\text{obs}})$ $=E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)\times\text{CDF}\left[\chi^{2}_{0}(T_{\text{obs}}/\tau_{\text{obs}}(E_{\gamma},z))\right]$
$\approx E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)\times\sqrt{\frac{2}{\pi}\frac{T_{\text{obs}}}{\tau_{\text{obs}}(E_{\gamma},z)}},\text{if }T_{\text{obs}}\gg\tau_{\text{obs}}$
@@
$\approx 0.36\times(1+z)^{(4+\alpha/2)}\left(\frac{B_{0}}{10^{-16}\text{ G}}\right)^{-1}\left(\frac{E_{\gamma}}{100\text{ GeV}}\right)$
×$\;\;\;\;\;\;\times E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)$ (2.9)
@@
where $dN/dE_{\gamma}(T_{\text{obs}}\rightarrow\infty)$ is the cascade spectrum of a steady-state source normalized to the same luminosity. The second line is true only in the limit where the observation time $T_{\text{obs}}$ is much smaller than $\tau_{\text{obs}}$. The last line is a rough scaling that is valid, e.g. when observing a source that is about 10 Mpc away for $T_{\text{obs}}=1~{}\text{yr}$ if $E_{\gamma}\ll\mathcal{O}(10\text{ TeV})$ and $B_{0}\gg\mathcal{O}(10^{-14}\text{ G})$. We compute the inclusive cascade spectra using $\gamma$-Cascade and predict the finite-observation spectrum using our simplified model. We note that since $\theta_{obs}\ll 1^{\text{o}}$ for observation times $T_{\text{obs}}\sim\mathcal{O}(\text{yrs})$, the angular extension of the cascade will not be resolved within the PSF of Fermi-LAT. Therefore, the only transient observable we consider is the afterglow “echo” of the electromagnetic cascades.
@@
Figure 1: Comparison of visible cascades for $T_{\text{obs}}=1$ year, $z=0.1$, and varying IGMF strengths. Our results, shown in black, are compared to those of Ref. [23], shown in red. The black curves correspond to computations performed using $\gamma$-Cascade, while the red curves are derived from ELMAG, a Monte Carlo simulation. The green dashed line represents the approximation given in Eq. 2.9, applied to the universal spectrum of a fully developed electromagnetic cascade, $dN/dE_{\gamma}\propto E_{\gamma}^{-1.9}$ [19], from which we predict an energy spectrum as a power-law with spectral index of $\Gamma\sim$ 0.9. The blue curve depicts the expected KM3NeT neutrino spectrum under the assumption of photo-pion production.

### v2

@@ 1 Introduction
The unexpected detection of a signal by the KM3NeT/ARCA Collaboration that is consistent with a $\gtrsim$100 PeV neutrino has potentially opened a new window into the most extreme accelerators in the Universe [1]. If this event is confirmed to be an astrophysical neutrino, KM3-230213A stands out as the highest energy neutrino observed to date, raising questions regarding its astrophysical [2, 3, 4, 5] or cosmological [6, 7] origins, as well as potential multiwavelength counterparts. This event prompted a variety of theoretical interpretations involving beyond Standard Model physics [8], including Lorentz invariance violation [9, 10, 11] and dark matter decay or interactions [12, 13, 14, 15, 16, 17, 18]. Scenarios involving evaporating primordial black holes have also been explored [19, 20]. The tension between the KM3NeT signal and the lack of similar detections by the larger IceCube detector motivates scenarios where the KM3NeT signal is produced by a transient source fortuitously located in the optimal region for KM3NeT detection [21, 22].
If the neutrino flux is generated via the hadronic interactions of ultra-high energy protons or nuclei, there should also be an accompanying flux of high-energy $\gamma$-rays [23, 24, 25, 26]. These $\gamma$-rays can initiate electromagnetic cascades on the extragalactic background light (EBL), shifting the bulk of their energy into the GeV–TeV band, where they could be observable by the Fermi Large Area Telescope (LAT). In the case of steady-state sources, these cascades could produce a diffuse $\gamma$-ray halo surrounding the source, the size and spectrum of which would depend sensitively on the intergalactic magnetic field (IGMF) strength. In the case of transient sources, the spectral evolution of the cascade is more difficult to model; however, the spatial evolution is simpler, as the $\gamma$-ray cascade produced over short time delays has a spatial extension much smaller than the resolution of current $\gamma$-ray instruments.
@@ 2.1 Electromagnetic Cascade Modeling
We model the electromagnetic counterpart of the neutrino signal under the assumption that a monoenergetic hadronic transient event at energies of $\mathcal{O}(100\text{ PeV})$ gave rise to the detected neutrino. We set the source luminosity to the total neutrino energy flux as calculated by KM3NeT, $E_{\nu}^{2}\Phi_{\nu}=5.8\times 10^{-8}\text{ GeV}\;\text{cm}^{-2}\;\text{s}^{-1}\;\text{sr}^{-1}$, which allows us to normalize the injected $\gamma$-ray energy up to a constant that accounts for the uncertainty in the production mechanism. We adopt a model for the injected $\gamma$-ray luminosity as follows:
@@
where $\eta\sim\mathcal{O}(1)$ accounts for the uncertainty in the production mechanism (e.g. $\eta=1$ for photo-pion production or $\eta=1/2$ for proton-proton interactions), and $dN(z)/dE_{\nu}dA=4\pi\Phi_{\nu}\tau_{\text{ARCA}}$ is the integrated differential number flux of neutrinos detected over an active observation time of KM3NeT/ARCA high-energy neutrino detector, $\tau_{\text{ARCA}}\approx 1\;\text{year}$. Furthermore, we assume that the $\gamma$-rays are produced with a spectral shape that is centered about 440 PeV and whose width is narrow enough such that essentially all the injected energy is reprocessed into cascade photons. Practically, we model this as a log-normal distribution with a variance (width) of up to a decade. We note that our results are not sensitive to the spectral width of the injected $\gamma$-rays as long as this width is less than $\sim$3 decades. At larger widths, there would be a significant number of unattenuated TeV $\gamma$-rays for distances below $\sim$10 Mpc. If the emission extends below 100 GeV, there would be an additional unattenuated $\gamma$-ray component which may be bright compared to the KM3NeT flux, and would typically strengthen the resulting constraints in this paper.
@@
All the primary $\gamma$-rays in our model will be attenuated due to pair production with the background radiation field stemming from a combination of the cosmic microwave background (CMB) and the EBL. We adopt the EBL model from Saldana-Lopez et al. (2021) [29]. After $e^{+/-}$ (hereafter, collectively called electrons) pair produce, these electrons are able to inverse-Compton scatter with the same background radiation, producing $\gamma$-rays. This cascade cycle reprocesses the energy above $\sim\mathcal{O}(\text{TeV})$ into GeV photons that are visible to the Fermi-LAT. We note that while the highest-energy electrons experience Klein-Nishina (KN) suppression of their cooling rates, the cascade quickly develops a population of lower-energy electrons ($\sim$1–100 TeV) for which Thomson scattering provides an accurate description of the energy losses [30]. We compute these cascades semi-analytically using $\gamma$-Cascade [27]. Figure 1 shows a comparison of our cascade spectra with Monte Carlo results from the ELMAG package, demonstrating good agreement between the two approaches for different IGMF strengths.
@@
As the primary $\gamma$-rays leave the source at comoving distance $d_{s}$, they will cascade and change their original direction of travel. This is caused by the motion of the electrons under the influence of the IGMF. Once electrons are generated via pair production, they experience an acceleration perpendicular to their velocity, which causes them to turn with a Larmor radius, $R_{L}=E_{e}/eB(z)$, where $B(z)=B_{0}(1+z)^{-\alpha}$ is the IGMF strength at redshift $z$. Here we take $\alpha=4$ for IGMF evolution driven by a dynamo mechanism, as predicted by theoretical models (see e.g. [31, 32, 33]). This treatment assumes that the IGMF coherence length $\lambda_{B}$ is much larger than the electron cooling length, allowing us to define a meaningful Larmor radius over the cascade development scale. This assumption is well-justified for typical IGMF parameters: we adopt coherence lengths $\lambda_{B}\sim 1$ Mpc, while inverse Compton cooling against the CMB occurs over much shorter distances of $\sim$10 kpc for 1 PeV electrons (including KN suppression) and $\sim$40 kpc for 10 TeV electrons in the Thomson regime. At 440 PeV, the KN-corrected cooling length is $\sim 0.3$ Mpc. Thus, $\lambda_{B}>d_{\rm cool}$ even at the earliest, KN-suppressed stage, and $\lambda_{B}\gg d_{\text{cool}}$ once the cascade develops into the TeV range that dominates the final deflection [30]. The total angle by which the electrons’ trajectory becomes deflected is given by,
$\delta=c\int_{E_{e,0}}^{E_{e}}dE_{e}^{\prime}\left(\frac{dE^{\prime}_{e}}{dt}\right)^{-1}\frac{1}{R_{L}},$ (2.2)
@@
and where $u_{\text{CMB}}$ is the energy density of the background radiation field, $\sigma_{T}$ is the Thomson cross section, and $\gamma_{e}=E_{e}/m_{e}c^{2}$ is the gamma factor of the electrons. Note that while we show the Thomson limit in Eq. 2.3 for simplicity, our $\gamma$-Cascade simulations include full Klein-Nishina corrections for inverse Compton scattering at all energies. We can evaluate the integral above in terms of the energy-loss distance of the electrons, $d_{\text{ICS}}=cE_{e}/(dE_{e}/dt)$,
$\delta(E_{e},z)$ $=c\int_{E_{e,0}}^{E_{e}}dE_{e}^{\prime}\left(\frac{dE^{\prime}_{e}}{dt}\right)^{-1}\frac{1}{R_{L}}\approx\frac{d_{\text{ICS}}(E_{e})}{2R_{L}(E_{e})}$
$=\frac{3m_{e}^{2}c^{4}eB(z)}{8\sigma_{T}u_{\text{CMB}}(z)E_{e}^{2}}$
$=\frac{3m_{e}^{2}c^{4}eB_{0}}{8\sigma_{T}u_{\text{CMB},0}E_{e}^{2}(1+z)^{4+\alpha}}.$ (2.4)
@@
We further assume a small-angle approximation, where $\delta\ll\pi$, (i.e. $d_{\text{ICS}}\ll R_{L}$). As the cascade progresses, the mean free path of the photons grows rapidly below $E_{\gamma}\sim\mathcal{O}(\text{PeV})$. At energies above the point where the mean free path reaches its minimum, the mean free path scales as $D_{\gamma\gamma}=\langle\bar{\sigma}_{\gamma\gamma}(E_{\gamma})n_{\text{CMB}}\rangle^{-1}\propto E_{\gamma}$. Therefore, one can think of the cascade as being composed of several cycles that extend out from the source up to characteristic distances set by the mean free path of the second-to-last pair-production step, $\ell_{\rm pair}(z)\sim D_{\gamma\gamma}(E_{\gamma}^{0},z)$, where $E_{\gamma}^{0}$ is the energy of that second-to-last photon (and the final-generation electron has energy $E_{e}=E_{\gamma}^{0}/2)$. After this point, the original direction of the cascading photons has been changed by $\delta$ from it’s original trajectory.
@@
The “last” photon is the one that is detected on Earth by the observer coming from a direction $\theta_{\text{obs}}$ with respect to the line of sight to the source. The observed photon energy $E_{\gamma}$ from the final inverse-Compton scattering on the CMB is related to the last-generation electron energy by $E_{\gamma}\simeq 4/3\left(E_{e}/m_{e}c^{2}\right)^{2}\epsilon_{\rm CMB}$, implying $E_{e}\propto\sqrt{E_{\gamma}}$ [34] [^1] This scaling underlies the $E_{\gamma}^{-1}$ dependence of $\theta_{\text{obs}}$ in Eq. 2.6. It has been shown in several studies that this angle, which manifests in the characteristic apparent angular extension of the source, can be approximated by the following,
$\theta_{\text{obs}}(E_{\gamma},z)$ $\approx\frac{D_{\gamma\gamma}({E_{\gamma}^{0}},z)}{d_{A}}\delta(E_{e}=E_{\gamma}^{0}/2,z),\text{and}$ (2.5)
@@
$\theta_{\text{obs}}(E_{\gamma},z=0)$ $\approx 1^{\prime\prime}\times\left(\frac{E_{\gamma}}{100\;\text{GeV}}\right)^{-1}\left(\frac{B_{0}}{10^{-16}\;\text{G}}\right)\left(\frac{D_{\gamma\gamma}(E^{0}_{\gamma},z)/d_{A}}{100}\right)^{-1},$ (2.6)
@@
where $d_{A}=d_{c}/(1+z)$ is the angular diameter distance. We approximate the profile of the photons arriving at earth by a Gaussian distribution $\mathcal{N}(\mu_{\theta}=0,\sigma_{\theta}=\theta_{\text{obs}})$ centered at $\mu_{\theta}=0$ with a characteristic width of $\sigma_{\theta}=\theta_{\text{obs}}(E_{\gamma},z)$. For a given observed angle, the observed time delay due to the additional travel distance of the cascade is given by,
$\tau_{\text{obs}}(E_{\gamma},z)$ $\approx(1+z)\frac{D_{\gamma\gamma}(E_{\gamma}^{0},z)}{2}\delta^{2}(E_{\gamma}^{0}/2,z)$
@@
$\approx 5\,\text{yrs}\times(1+z)^{-(7+2\alpha)}\left(\frac{D_{\gamma\gamma}(E_{\gamma}^{0},z)}{10\;\text{Mpc}}\right)\left(\frac{B_{0}}{10^{-16}\;\text{G}}\right)^{2}\left(\frac{E_{\gamma}}{100\;\text{GeV}}\right)^{-2}.$ (2.7)
@@
Notice that since $\tau_{\text{obs}}\propto\theta_{\text{obs}}^{2}$, the time delay of the $\gamma$-rays is distributed as a chi-squared distribution with one degree of freedom, $\chi^{2}_{1}((\tau-\tau_{s})/\tau_{\text{obs}})$, where $\tau_{s}$ is the light-travel time to the source and $\tau$ is the time at which an observation is made . This is a consequence of our simplified model where the photons’ arrival is distributed as an angular Gaussian. A key result is that the total integrated flux of $\gamma$-rays collected after a total observation time $T_{\text{obs}}$ is given by the CDF of this chi-squared distribution,
$E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},T_{\text{obs}})$ $=E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)\times\text{CDF}\left[\chi^{2}_{1}(T_{\text{obs}}/\tau_{\text{obs}}(E_{\gamma},z))\right]$
$\approx E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)\times\sqrt{\frac{2}{\pi}\frac{T_{\text{obs}}}{\tau_{\text{obs}}(E_{\gamma},z)}},\text{if }T_{\text{obs}}\ll\tau_{\text{obs}}$
@@
$\approx 0.36\times(1+z)^{(7/2+\alpha)}\left(\frac{B_{0}}{10^{-16}\text{ G}}\right)^{-1}\left(\frac{E_{\gamma}}{100\text{ GeV}}\right)$
$\;\;\;\;\;\;\times E_{\gamma}^{2}\frac{dN}{dE_{\gamma}}(z,E_{\gamma},\infty)$ (2.8)
@@
where $dN/dE_{\gamma}(T_{\text{obs}}\rightarrow\infty)$ is the cascade spectrum of a steady-state source normalized to the same luminosity. The second line is true only in the limit where the observation time $T_{\text{obs}}$ is much smaller than $\tau_{\text{obs}}$. The last line is a rough scaling that is valid, e.g. when observing a source that is about 10 Mpc away for $T_{\text{obs}}=1~\text{yr}$ if $E_{\gamma}\ll\mathcal{O}(10\text{ TeV})$ and $B_{0}\ll\mathcal{O}(10^{-14}\text{ G})$. We compute the inclusive cascade spectra using $\gamma$-Cascade and predict the finite-observation spectrum using our simplified model. We note that since $\theta_{obs}\ll 1^{\text{o}}$ for observation times $T_{\text{obs}}\sim\mathcal{O}(\text{yrs})$, the angular extension of the cascade will not be resolved within the PSF of Fermi-LAT. Therefore, the only transient observable we consider is the afterglow “echo” of the electromagnetic cascades.
@@
Figure 1: Comparison of visible cascades for $T_{\text{obs}}=1$ year, $z=0.1$, and varying IGMF strengths, computed assuming IGMF evolution with $\alpha=4$. Our results, computed using the semi-analytic $\gamma$-CascadeV4 code, are shown in darker shades of teal-green curves, and are compared to the Monte Carlo-based ELMAG package in Ref. [34] shown in lighter shades. The orange dashed line represents the approximation given in Eq. 2.8, applied to the universal spectrum of a fully developed electromagnetic cascade, $dN/dE_{\gamma}\propto E_{\gamma}^{-1.9}$ [26], from which we predict an energy spectrum as a power-law with spectral index of $\Gamma\sim$ 0.9. The expected KM3NeT neutrino spectrum under the assumption of photo-pion production is shown in pink.